Actions
June 24th, 2015 (Evening Run)¶
Aim: Finish diffuser tests, high/low/nominal reference momentum pion beams for tracker alignment
Daily schedule¶
June | 24th | |
Time | 10pm - 6am | |
MOM | R. Bayes | |
Lead shifter | C. Pidcott | |
Shifter 2 | A. Dick | |
BLOC | C. Pidcott | |
SOC | J. Nugent | |
Experimenter | ||
Physics Shifter | P. Franchini | |
Beam? | Yes | |
Additional Personnel |
Shift Monitoring¶
Shifters should keep a careful eye on:- Water chiller temperature outside the control room (bottom of the red LEDs). If this reaches 25C, immediately power down the magnets and inform the MOM.
- Magnet water temperatures.
- Log any abnormal changes in the eLog
- If yellow temperature alarm sounds, immediately power down the magnets and inform the MOM.
- Magnet water flows.
- Log any abnormal changes in the eLog
- If yellow alarm sounds, immediately power down the magnets and inform the MOM.
- EMR stopping the DAQ arming
- Power cycle the VME crate by following the instructions here
Shift plan:¶
- Post Daily Meeting
- Shifter controlled entry training prior to shift
- Henry to do a quad soak test to verify upper limit to quad currents.
- 8:45pm
- Shifters and BLOC arrive to complete Hall preparation (i.e. target prep)
- 10pm
- Complete diffuser studies continuing from where they were left off on Monday.
- When complete, take 1/3 of shift with the 190 MeV pion beam (settings below)
- When complete, take 1/3 of shift with the 300 MeV pion reference beam (i.e. in run control) if not contra-indicated.
- When complete, take 1/3 of shift with the 290 MeV pion beam (settings below)
- Check that POMPOMs read magnet polarity correctly for input to CDB
- Photograph/Screenshot Run Summary at end of run for comparison to CDB (see screenshot attached to https://micewww.pp.rl.ac.uk/elog/MICE+Log/3223)
- Write more verbose eLogs, describe the problem, its symptoms and solution (if known).
Data Requested¶
- Pion reference beam for diffuser studies.
- 1/3 remainder of shift with 190 MeV pion beam, 1/3 shift on the 290 MeV pion beam, and 1/3 with the nominal beam.
Beamline Configuration¶
- We expect to run with 2V beamloss and the beam bump active.
- DS is off
- Beam settings via Run Control and as given below
Pion Beam for 190 MeV/c without Decay Solenoid¶
Number of particle triggers | Particle Species | p at Tgt | p@D1 | p@D2 | p@Tof0 | Proton Absorber | Q1 | Q2 | Q3 | D1 | DS | D2 | Q4 | Q5 | Q6 | Q7 | Q8 | Q9 | Similar Run(s) | RunControl tag |
# | MeV/c | MeV/c | MeV/c | MeV/c | mm | A | A | A | A | A | A | A | A | A | A | A | A | DAQ run # | ||
pion | 226.98 | 222.83 | 219.82 | 29 | 47.95 | 87.63 | 53.61 | 167.48 | - | 83.66 | 145.81 | 195.54 | 129.56 | 125.01 | 189.03 | 161.27 | LowRefPionRun_noDS |
Pion Beam for 290 MeV/c without Decay Solenoid¶
Number of particle triggers | Particle Species | p at Tgt | p@D1 | p@D2 | p@Tof0 | Proton Absorber | Q1 | Q2 | Q3 | D1 | DS | D2 | Q4 | Q5 | Q6 | Q7 | Q8 | Q9 | Similar Run(s) | RunControl tag |
# | MeV/c | MeV/c | MeV/c | MeV/c | mm | A | A | A | A | A | A | A | A | A | A | A | A | DAQ run # | ||
pion | 319.97 | 316.45 | 313.94 | 29 | 67.99 | 124.32 | 76.09 | 241.68 | - | 117.80 | 208.99 | 280.27 | 185.91 | 190.29 | 288.04 | 246.14 | HighRefPionRun_noDS |
Phone details for 6pm daily meeting¶
Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402
Updated by Bayes, Ryan almost 8 years ago ยท 1 revisions